OpenStudy (anonymous):

-4x+y=7 and 12x-3y=3 need to use substitution method to find x and y

OpenStudy (texaschic101):

Let me get you started.. -4x + y = 7 --- add 4x to both sides y = 4x + 7 12x - 3y = 3 -- sub in 4x + 7 in for y 12x - 3(4x + 7) = 3 Now just distribute through the parenthesis and solve for x. Once x is found, sub that back into either of the original equations and solve for y